# GAIA: A Global, Multi-modal, Multi-scale Vision-Language Dataset for Remote Sensing Image Analysis

### v1

@@ II-C The imperative need for better annotations
Subsequent investigations have explored a variety of techniques aimed at refining and augmenting existing datasets. One prominent approach involves improving the caliber of captions through the utilization of sophisticated language models. For instance, Fan et al. [62] proposed a method to improve the quality of captions by leveraging language rewrites. They used a large language model, specifically GPT-3, to rewrite existing captions in the Conceptual Captions and SBU Captions datasets, focusing on improving the descriptiveness and accuracy of the text. This method was shown to improve the performance of CLIP models on downstream tasks, demonstrating the importance of high-quality captions for contrastive language-image pre-training.
Concurrently, Nguyen et al. [63] focused on refining existing captions in multimodal datasets to improve the training of multimodal models. They utilized BLIP-2, a state-of-the-art vision-language model, to generate synthetic captions, and found that incorporating these captions during training led to significant performance gains in downstream tasks, further highlighting the importance of caption quality.
@@
In a paradigm shift, Yu et al. [65] introduced CapsFusion, a novel approach that rethinks image-text data at scale. CapsFusion leverages an ensemble of captioning models, including BLIP, BLIP-2, and GIT, to generate a diverse set of captions for each image. These captions are then fused using a large language model, specifically Vicuna-7B, to create a more comprehensive description. This method addresses the limitations of single-model captioning and demonstrates the advantages of leveraging multiple models to enhance caption quality.
@@
Xu et al. [66] investigated the potential of alt-text as a valuable source of image descriptions. Their work, altogether, introduced a method for re-aligning alt-text with images, thereby leveraging this often underutilized information source. They employed a multi-stage training process involving human annotation and a specifically designed captioning model to enhance the alignment between alt-text and image content. This method was shown to improve image captioning performance, highlighting the value of alt-text in providing rich visual descriptions.
@@
The application of synthetic data was further explored in specialized domains, such as medical vision-language pre-training. Liu et al. [67] investigated the feasibility of training effective medical vision-language models using purely synthetic data. They created a synthetic dataset of chest X-ray images and reports using generative models and demonstrated that models trained on this synthetic data could achieve comparable performance to those trained on real data, thus addressing data scarcity and privacy concerns inherent in medical datasets.
@@
Rotstein et al. [68] presented FuseCap, a system that leverages large language models to generate enriched, fused image captions. FuseCap employs frozen vision encoders, including an object detector, an attribute recognizer, and an OCR model, to extract detailed visual information from images. These outputs are then fused with original captions using a fine-tuned Flan-T5 model, resulting in comprehensive image descriptions. They curated a dataset of 12 million image-enriched caption pairs and demonstrated that training a BLIP-based captioning model on this data led to improved performance on downstream tasks.
@@
Advancing the use of synthetic data, Sharifzadeh et al. [69] introduced a method that boosts visual-language models by utilizing synthetic captions and image embeddings. This approach employs a text-to-image model (MUSE) to generate synthetic image embeddings from captions generated by a large language model (LLaMA-3). They demonstrated that fine-tuning a VLM on this synthetic data achieves comparable performance to models trained on real data, highlighting the efficacy of combining different types of synthetic data.
@@
Finally, Li et al. [70] undertook a large-scale experiment to investigate the capabilities of state-of-the-art language models in generating high-quality captions. They employed a LLaMA-3-powered LLaVA model to recaption 1.3 billion web images from the DataComp-1B dataset. Their results demonstrated significant improvements in the quality of the generated captions compared to the original web-crawled captions, highlighting the potential of advanced language models in large-scale data augmentation.
@@ III-A1 Data Collection
Source Original Filtered % Retained
NASA 54,577 36,239 66.4
ESA 5,932 3,528 59.5
NOAA 651 455 69.9
AIRBUS 413 409 99.0
Planet 612 399 65.2
Total 62,185 41,030 66.0
@@ III-A2 Data Cleaning and De-duplication
The raw web-scraped dataset underwent a data cleaning and de-duplication phase. Textual data was cleaned by removing HTML tags and URLs. To address instances of empty alt-text HTML tags, the corresponding article’s title was stored as a substitute of the original image caption. Image de-duplication was performed in two distinct phases. Initially, duplicate images were identified based on their source URLs. Subsequently, we utilized fastdup  [^3] https://github.com/visual-layer/fastdup, an open-source tool for detecting visually similar images, to identify more duplicates, outliers, and corrupted images. In cases of duplication, dataset entries were merged, including their metadata, while outliers and corrupted images were eliminated from the dataset.
@@ III-A3 Data Annotation Pipeline
To facilitate the extraction of relevant metadata and the generation of comprehensive captions, a detailed prompt was developed to instruct ChatGPT and specifically their most capable model GPT-4o. The model was tasked with analyzing each RS image and its associated article’s text, extracting relevant domain-specific metadata, and generating five distinct, detailed captions per image. The prompt employed a system of rewards and penalties to reinforce desirable and discourage undesirable caption characteristics, respectively. Furthermore, a structured output schema was integrated to ensure consistent formatting and prevent omission of required fields or inclusion of hallucinated information. This structured approach promotes the extraction of domain-specific metadata and captions, enhancing consistency and facilitating downstream tasks such as dataset querying, filtering, and analysis. The systematic combination of detailed prompting and structured output contributed to the creation of a rich and informative dataset tailored for EO and RS applications. An overview of the extracted metadata fields and their descriptions is presented in Table III. For excessively long articles, a text summarization step was prepended to the data annotation pipeline. This involved utilizing the far more cost-effective, yet capable for the summarization task, GPT-4o-mini model to generate concise summaries, thereby mitigating annotation costs for the following annotation pipeline step while preserving the quality of extracted metadata and captions. Our qualitative analysis revealed that the geographic coordinates (EPSG:4326) generated by GPT-4o, while generally proximate, lacked the desired precision. Consequently, the GPT-4o generated coordinates were discarded in favor of those obtained from the Google Geocoding API  [^4] https://developers.google.com/maps/documentation/geocoding/requests-geocoding, which provides accurate latitude and longitude coordinates through toponym resolution.
@@ III-B Exploratory Data Analysis
GAIA’s multi-modal nature is highlighted in Fig. 7, which displays the distribution of RS modalities. While optical/multispectral imagery constitutes the majority (88.1%), the dataset also incorporates thermal infrared (6.6%), radar (2.6%), microwave radiometry (1.0%), and near-infrared (0.5%) data, among others.
@@
   Size 16,173 28,751
   Hapax Legomena 6,543 7,862
   Total Words 394,471 7,493,402
@@
Focusing on the textual characteristics of the GAIA dataset, a comparative analysis between the original alt-text captions and our synthetic captions, shown in Table IV, reveals significant differences in their richness and complexity. The synthetic captions exhibit a considerably larger vocabulary (28,751 vs. 16,173) and a higher number of unique words (Hapax Legomena of 7,862 vs. 6,543), indicating a more diverse and nuanced use of language. This enhanced lexical diversity is further emphasized by the substantially higher total word count in the synthetic captions (7,493,402 vs. 394,471). Furthermore, the average number of words per caption is significantly greater for the synthetic (36.53) compared to the original alt-text (9.61), with a lower standard deviation (7.23 vs. 12.48), suggesting not only longer but also more consistently detailed descriptions. Each image is associated with five synthetic captions, compared to a single alt-text caption, resulting in a fivefold increase in the total number of captions and offering a diverse textual representation of each scene.
@@ IV-B Experimental Results
A qualitative analysis of the generated captions (see Fig. 12) provides deeper insights into the fine-tuned model’s strengths and weaknesses, which might be traced back to GAIA and reveal potential limitations or potent areas for improvement. Fine-tuned models, particularly those trained on synthetic captions, generate predictions that are more relevant to the image content and reference captions compared to zero-shot predictions. They are better at identifying key elements, locations, and events described in the references captions. Moreover, GTP-4o fine-tuned models show improved ability to handle complex RS scenes, such as those involving natural disasters (wildfires, floods, cyclones) and environmental phenomena (phytoplankton blooms, dust storms). They more accurately identify these events compared to their other two counterparts. Last but not least, models fine-tuned using our synthetic captions generate prominently more extensive and informative descriptions.
@@
Alt-Text Zero-shot $0.00$ $14.05$ $3.27$ $13.32$ $12.35$
Fine-tuned $6.34$ $26.74$ $13.30$ $26.32$ $26.44$
Synthetic Zero-shot $0.26$ $22.19$ $5.77$ $18.29$ $9.35$
Fine-tuned $25.76$ $42.64$ $19.64$ $33.91$ $36.90$

### v2

@@ II-C The imperative need for better annotations
Subsequent investigations have explored a variety of techniques aimed at refining and augmenting existing datasets. One prominent approach involves improving the caliber of captions through the utilization of sophisticated language models. Leveraging language rewrites has been shown to improve the quality of captions [22]. In one such approach, a large language model (GPT-3) was used to rewrite captions from the Conceptual Captions and SBU Captions datasets to enhance their descriptiveness and accuracy. This method was shown to improve the performance of CLIP models on downstream tasks, demonstrating the importance of high-quality captions for contrastive language-image pre-training.
Concurrently, related work has focused on refining captions in multi-modal datasets to improve model training [61]. In this approach, a state-of-the-art vision-language model (BLIP-2) was used to generate synthetic captions. Incorporating these new captions during training was found to lead to significant performance gains in downstream tasks, further highlighting the importance of caption quality.
@@
In a paradigm shift, CapsFusion [91] demonstrated a novel approach that rethinks image-text data at scale. CapsFusion leverages an ensemble of captioning models, including BLIP, BLIP-2, and GIT, to generate a diverse set of captions for each image. These captions are then fused using a large language model, specifically Vicuna-7B, to create a more comprehensive description. This method addresses the limitations of single-model captioning and demonstrates the advantages of leveraging multiple models to enhance caption quality.
@@
The potential of alt-text as a valuable source of image descriptions, has also been investigated [88], introducing a method for re-aligning alt-text with images and thereby leveraging this often underutilized information source. They employed a multi-stage training process involving human annotation and a specifically designed captioning model to enhance the alignment between alt-text and image content. This method was shown to improve image captioning performance, highlighting the value of alt-text in providing rich visual descriptions.
@@
The application of synthetic data was further explored in specialized domains, such as medical vision-language pre-training [49]. They investigated the feasibility of training effective medical vision-language models using purely synthetic data. They created a synthetic dataset of chest X-ray images and reports using generative models and demonstrated that models trained on this synthetic data could achieve comparable performance to those trained on real data, thus addressing data scarcity and privacy concerns inherent in medical datasets.
@@
FuseCap [69] introduced a system that leverages large language models to generate enriched, fused image captions. FuseCap employs frozen vision encoders, including an object detector, an attribute recognizer, and an OCR model, to extract detailed visual information from images. These outputs are then fused with original captions using a fine-tuned Flan-T5 model, resulting in comprehensive image descriptions. They curated a dataset of 12 million image-enriched caption pairs and demonstrated that training a BLIP-based captioning model on this data led to improved performance on downstream tasks.
@@
Advancing the use of synthetic data, a method that boosts visual-language models by utilizing synthetic captions and image embeddings was introduced [74]. This approach employs a text-to-image model (MUSE) to generate synthetic image embeddings from captions generated by a large language model (LLaMA-3). They demonstrated that fine-tuning a VLM on this synthetic data achieves comparable performance to models trained on real data, highlighting the efficacy of combining different types of synthetic data.
@@
Finally, a large-scale experiment to investigate the capabilities of state-of-the-art language models in generating high-quality captions has been conducted [43]. A LLaMA-3-powered LLaVA model was employed to recaption 1.3 billion web images from the DataComp-1B dataset. Their results demonstrated significant improvements in the quality of the generated captions compared to the original web-crawled captions, highlighting the potential of advanced language models in large-scale data augmentation.
@@ III-A1 Data Collection
Source Domain(s) Original Filtered % Retained
NASA visibleearth.nasa.gov 54,577 35,468 65.0
earthobservatory.nasa.gov
ESA esa.int 5,932 3,487 58.8
copernicus.eu
NOAA nesdis.noaa.gov 651 438 67.3
AIRBUS space-solutions.airbus.com 413 409 99.0
Planet planet.com 612 399 65.2
Total 62,185 40,201 66.0
@@ III-A2 Data Cleaning and De-duplication
The raw web-scraped dataset underwent a data cleaning and de-duplication phase. Textual data was cleaned by removing HTML tags and URLs. To address instances of empty alt-text HTML tags, the corresponding article’s title was stored as a substitute of the original image caption. Image de-duplication was performed in three distinct phases. Initially, duplicate images were identified based on their source URLs. Subsequently, we utilized fastdup  [^4] https://github.com/visual-layer/fastdup, an open-source tool to identify more duplicates, outliers, and corrupted images. In cases of duplication, dataset entries were merged, including their metadata, while outliers and corrupted images were eliminated from the dataset. Ultimately, we implemented a multi-stage filtering protocol to further enhance dataset quality by identifying and removing remaining outliers and low-quality samples. For this, we leveraged a state-of-the-art vision-language model, PE-Core-G14-448 [6], to score each image based on its semantic similarity to a set of text prompts (e.g., ’a satellite image’, ’an earth observation image’, etc.). Our filtering process proceeded in three stages. First,  30 images with extremely low similarity scores were flagged; manual inspection confirmed these were remaining outliers, and they were subsequently removed. Second, to address images of insufficient quality, we manually evaluated the next 4,000 lowest-scoring samples, representing approximately 10% of the dataset. This rigorous review led to the removal of  800 images that, while relevant, failed to meet our quality standards. Criteria for removal included poor visual fidelity (e.g., low-resolution or blurry images), the presence of rendered content, and scientific visualizations that lacked essential interpretative elements like legends or labels. Finally, to validate the integrity of the resulting dataset, we performed a final quality check on a random subset of 1,000 images. This final inspection confirmed the effectiveness of our data cleaning and de-duplication protocol, as no further issues were identified.
@@ III-A3 Data Annotation Pipeline
To facilitate the extraction of relevant metadata and the generation of comprehensive captions, a detailed prompt was developed to instruct ChatGPT and specifically their most capable model GPT-4o (gpt-4o-2024-08-06). The model was tasked with analyzing each RS image and its associated article’s text, extracting relevant domain-specific metadata, and generating five distinct, detailed captions per image. Following OpenAI’s best practices guide, we defined effective instructions for the model, such that it consistently generates content that meets our requirements and a structured output schema that ensures the model will always generate responses that adhere to our supplied JSON schema, without omitting or hallucinating required fields. The prompt employed a system of rewards and penalties to reinforce desirable and discourage undesirable caption characteristics, respectively. This structured approach promotes the extraction of domain-specific metadata and captions, enhancing consistency and facilitating downstream tasks such as dataset querying, filtering, and analysis. The systematic combination of detailed prompting and structured output contributed to the creation of a rich and informative dataset tailored for EO and RS applications. An overview of the extracted metadata fields and their descriptions is presented in Table III. For excessively long articles, a text summarization step was prepended to the data annotation pipeline. This involved utilizing the far more cost-effective, yet capable for the summarization task, GPT-4o-mini (gpt-4o-mini-2024-07-18) model to generate concise summaries, thereby mitigating annotation costs for the following annotation pipeline step while preserving the quality of extracted metadata and captions. Our qualitative analysis revealed that the geographic coordinates (EPSG:4326) generated by GPT-4o, while generally proximate, lacked the desired precision. Consequently, the GPT-4o generated coordinates were discarded in favor of those obtained from the Google Geocoding API  [^5] https://developers.google.com/maps/documentation/geocoding/requests-geocoding, which provides accurate latitude and longitude coordinates through toponym resolution.
@@ III-B Exploratory Data Analysis
GAIA’s multi-modal nature is highlighted in Fig. 8, which displays the distribution of RS modalities. While optical/multispectral imagery constitutes the majority (88.1%), the dataset also incorporates thermal infrared (6.6%), radar (2.6%), microwave radiometry (1.0%), and near-infrared (0.5%) data, among others. This observed distribution largely reflects the prevailing trends in RS data availability and historical usage patterns, where optical sensors have traditionally been more widely deployed and utilized, often due to factors such as the ease of acquisition and interpretability. Nevertheless, the non-optical modalities provide unique information that optical data does not capture, enabling a wider range of studies and thus enabling a more complete understanding of the Earth.
@@
   Size 15,788 28,396
   Hapax Legomena 6,403 7,770
   Total Words 379,000 7,346,969
@@
Focusing on the textual characteristics of the GAIA dataset, a comparative analysis between the original alt-text captions and our synthetic captions, shown in Table IV, reveals significant differences in their richness and complexity. The synthetic captions exhibit a considerably larger vocabulary (28,396 vs. 15,788) and a higher number of unique words (Hapax Legomena of 7,770 vs. 6,403), indicating a more diverse and nuanced use of language. This enhanced lexical diversity is further emphasized by the substantially higher total word count in the synthetic captions (7,346,969 vs. 379,000). Furthermore, the average number of words per caption is significantly greater for the synthetic (36.55) compared to the original alt-text (9.43), with a lower standard deviation (7.20 vs. 12.34), suggesting not only longer but also more consistently detailed descriptions. Each image is associated with five synthetic captions, compared to a single alt-text caption, resulting in a fivefold increase in the total number of captions and offering a diverse textual representation of each scene.
To quantitatively assess the diversity of the five synthetic captions accompanying each image in the GAIA dataset, we conducted a comprehensive analysis. This analysis aimed to ensure that the multiple captions provide varied lexical and semantic perspectives on the visual content, thereby adding significant value over a single caption. Semantic diversity was evaluated by generating sentence embeddings using Sentence-BERT (SBERT) and calculating pairwise cosine similarities within each five-caption set. This yielded a global mean similarity of 0.7153 (median: 0.7421, std: 0.1135), corresponding to an average semantic diversity score of 0.2847 across all 40,201 sets. Concurrently, lexical diversity was evaluated using the Measure of Textual Lexical Diversity (MTLD) and Hypergeometric Distribution D (HD-D) on the combined and lemmatized text of each caption set. This achieved a mean MTLD score of 74.66 (median: 72.71, std: 16.03) and a mean HD-D score of 0.7746 (median: 0.7750, std: 0.0328). Collectively, these metrics indicate that while the five captions are semantically anchored to the same visual content, they exhibit moderate semantic diversity and notable distinctiveness in their phrasing, focus, and vocabulary richness, thereby affirming their significant value in providing multifaceted textual representations for each image.
As a final sanity check we employed gpt-4o-mini (gpt-4o-mini-2024-07-18) to verify the relevance of our generated captions with respect to the accompanied image. To this end, we prompted the model to evaluate the generated captions relevance and completeness and rate them with a score between 0 and 10 with 0 being the lowest and 10 the highest score. We end up with an overall score of 9.32 (see Fig. 13) proving the efficacy of our annotation pipeline to produce concise, accurate and scientifically relevant captions. Manual inspection of a random subset of the lower-scoring samples revealed two primary contributing factors: (a) limited context, i.e., the web-scraped text associated with the images was relatively short, and (b) constraints in the instruction-following capabilities of the model used for caption generation. While currently implemented as a retrospective validation measure, this metric possesses significant utility as an active quality-control signal for future model development. Leveraging established methodologies in robust vision-language learning, we identify three mechanisms to leverage this signal:
  1. 1.
    Iterative Self-Refinement: Adopting the Self-Refine paradigm [58], this score functions as a reward signal within an dataset caption generation feedback loop. The model can be prompted to iteratively regenerate captions until a maximization threshold is met, ensuring outputs satisfy the strictest criteria for completeness.
  2. 2.
    Threshold-based Filtering: Validation via CLAIR [11] establishes LLM-based metrics as superior proxies for human judgment compared to n-gram statistics. Consequently, this score can function as a semantic gatekeeper, allowing for the pruning of relatively lower-scoring samples to curate a subset with maximized aggregate semantic fidelity.
  3. 3.
    Quality-Aware Learning: The evaluation score can serve as a reliability index for adaptive loss weighting [32]. By dynamically modulating the gradient contribution of each sample based on its alignment quality, the optimization process prioritizes learning from the most information-rich captions without discarding any data.
Figure 13: Frequency distribution of 40,201 quality ratings assigned by gpt-4o-mini based on caption relevance and completeness. The heavily left-skewed distribution, with a mean of 9.32 (std=0.46) and a median of 9.39, provides robust quantitative evidence for our annotation pipeline’s efficacy. The tight clustering of ratings at the highest end of the 10-point scale confirms our annotation pipeline’s capacity to consistently generate concise, accurate, and scientifically relevant captions.
@@ IV-B Experimental Results
For completeness, we compare both our alt-text and synthetic captions fine-tuned CLIP-ViT-L-14 models with the zero-shot CLIP model and other RS-domain fine-tuned CLIP models, i.e. SkyCLIP [84], RemoteCLIP [51], GeoRSCLIP [97], LAION_RS [84]. Despite using a dataset several times smaller than those employed to fine-tune the rest of the models considered and training for only a single epoch, GAIA delivers strong classification and retrieval performance.
Table IX reveals that for the classification task, GAIA (Synthetic) ranks among the top-performing models. On EuroSAT and ImageNet1K, our GAIA (Synthetic) model positions itself among the top-performing methods. For EuroSAT, it achieves an Acc1 of 72.46%, outperforming SkyCLIP (70.54%) and LAION_RS (71.33%), and is competitive with RS5M (74.52%). On ImageNet1K, its Acc1 of 75.49% is on par with SkyCLIP and the baseline CLIP, and approaches LAION_RS (75.61%), demonstrating strong generalization to common vision tasks. Conversely, on RESISC45, the performance of our GAIA-tuned models is closer to the zero-shot CLIP baseline. This discrepancy may be attributed to the GAIA’s comparatively limited inclusion of high-resolution aerial imagery, a characteristic of RESISC45, as discussed in our exploratory data analysis (see Section III-B). Nevertheless, the overall results underscore GAIA’s potential for efficient, effective and semantically rich fine-tuning, yielding competitive performance with significantly reduced data and computational requirements.
On the other hand, for cross-modal retrieval, in order to better demonstrate GAIA’s unique features through RS tailored benchmarks, we employed the thematic tags information associated with each GAIA sample. More specifically, each GAIA sample is complemented by three generic tags, which encompass thematic aspects such as earth sphere, topic, and location (e.g., [”Cryosphere”, ”Climate Change”, ”Arctic”] or [”Floods”, ”Sediment Transport”, ”Adriatic Sea”]). We deliberately chose to base our benchmarks on these tags rather than on the aggregated five major earth spheres and their subcategories, shown in Fig. 11. This choice was motivated by two key considerations:
  1. 1.
    Higher Semantic Variance: The thematic tags provide richer semantic information compared to the earth sphere aggregates (5) and its subcategories which are relatively coarse and limited in variability.
  2. 2.
    Relationship Structure: Due to the fact that benchmarks and metrics commonly used in the context of CLIP, rely on one-to-one relationships for evaluation, GAIA’s thematic tags present a more challenging and realistic setting to validate models capacity for semantic understanding and cross-modal alignment. More specifically, the tags form complex one-to-one matches with images (¿6600 different thematic tag combinations within our GAIA test split), whereas the earth spheres (5) and subcategories typically yield simpler one-to-many mappings.
Table X demonstrates that our GAIA-fine-tuned models, particularly GAIA (Synthetic), establish a new state-of-the-art across all recall metrics when evaluated using GAIA’s thematic tags. For text-to-image retrieval, the GAIA (Synthetic) model achieves a Recall@100 of 47.32%, outperforming the zero-shot CLIP baseline (38.07%) and the second-best model, LAION_RS (45.03%). The advantage is even more pronounced in image-to-text retrieval, where GAIA (Synthetic) reaches a Recall@100 of 51.79%, creating a large margin over both the baseline CLIP (36.97%) and the rest of RS-domain fine-tuned models. Finally, our GAIA (Alt-text) model also delivers very strong performance, exhibiting the second-best results in image-to-text retrieval and third-best results in text-to-image retrieval.
@@
A qualitative analysis of the generated captions (see Fig. 14) provides deeper insights into the fine-tuned model’s strengths and weaknesses, which might be traced back to GAIA and reveal potential limitations or potent areas for improvement. Fine-tuned models, particularly those trained on synthetic captions, generate predictions that are more relevant to the image content and reference captions compared to zero-shot predictions. They are better at identifying key elements, locations, and events described in the references captions. Moreover, synthetic caption fine-tuned models show improved ability to handle complex RS scenes, such as those involving natural disasters (wildfires, floods, cyclones) and environmental phenomena (phytoplankton blooms, dust storms). They more accurately identify these events compared to their other two counterparts. Last but not least, models fine-tuned using our synthetic captions generate prominently more extensive and informative descriptions.
@@
Lastly, we further evaluate our best-performing models - the BLIP2 and CLIP-ViT-L-14 variants fine-tuned on GAIA’s synthetic captions - on two additional RS image-text paired datasets: RSICap and GIT-10M. In the absence of standard evaluation splits for GIT-10M, we curated a stratified test set of 100k samples from its georeferenced partition. Adopting the GAIA spatial stratification protocol, we maximized global geographic diversity while simultaneously balancing the dataset across a Ground Sample Distance (GSD) spectrum of 0.5m to 128m to ensure scale invariance. The resulting subset exhibits significant linguistic variance, characterized by a mean caption length of 52 words (min=7, max=127). The results, presented in Table VIII, summarize the performance of our GAIA fine-tuned models against their zero-shot baselines. For the cross-modal retrieval task, our CLIP-ViT-L-14 model exhibits significant improvements. For both text-to-image and image-to-text recall, our fine-tuned model consistently outperforms the zero-shot baseline across all Recall@k metrics. As for image captioning the results demonstrate that the BLIP2 model fine-tuned on GAIA achieves substantial performance gains on both external datasets. On RSICap, the METEOR score doubles from 7.35 to 15.09, and the ROUGE-L score increases from 15.21 to 18.76. The improvement is even more pronounced on GIT-10M, where the METEOR score jumps from 6.83 to 16.62. Collectively, these results demonstrate that the knowledge encoded in the GAIA dataset equips models with robust, generalizable knowledge that significantly enhances performance on a diverse range of RS downstream tasks. However, it is important to contextualize these improvements by highlighting the significant annotation differences between GAIA’s captions and those of other RS image-text paired datasets (see Fig. 2). While the performance gains are clear, reference-based metrics may not fully capture the semantic quality of the generated captions, as they are being compared solely against reference captions from datasets exhibiting different annotation paradigms. Typically, fine-tuning induces a linguistic bias that degrades performance on datasets with different annotation style. However, our analysis attributes the observed gains to two countervailing mechanisms:
  1. 1.
    Visual Feature Adaptation: The zero-shot BLIP-2 baseline has been pre-trained primarily on natural, web-crawled imagery, resulting in a significant domain gap when processing satellite/aerial perspectives. Fine-tuning on GAIA allows the model to learn these overhead visual features and subsequently relate their semantic content to domain-specific vocabulary. This ability to correctly identify objects is a fundamental improvement that applies to any dataset, regardless of the caption style.
  2. 2.
    Compensating for Metric Penalty: Conventional evaluation metrics rely heavily on n-gram overlap and inherently penalize stylistic deviations from the ground truth. Since GAIA induces a linguistic shift distinct from the styles of RSICap or GIT-10M, a degradation in these metrics would be the theoretical expectation. The fact that quantitative scores increased suggests that the gain in semantic fidelity, i.e. the ability to correctly infer content and utilize precise terminology, is substantial enough to outweigh the penalty imposed by the stylistic mismatch.
Last but not least, we must highlight the stark differences in terms of both visual feature recognition competence and the level of description detail and domain-specific vocabulary use of the zero-shot BLIP2 model prediction, observed in Fig. 14.
@@
Alt-Text Zero-shot $0.30$ $14.06$ $2.95$ $12.94$ $10.94$
Fine-tuned $4.81$ $23.78$ $10.77$ $23.09$ $22.09$
Synthetic Zero-shot $0.23$ $21.85$ $5.51$ $17.96$ $9.15$
Fine-tuned $25.62$ $41.48$ $19.30$ $33.35$ $36.08$
TABLE VIII: Remote Sensing Image Captioning (BLIP2) and Cross-Modal Retrieval Results (CLIP-ViT-L-14) on external datasets RSICap and GIT-10M
Dataset Mode Evaluation Metrics (%)
BLEU ROUGE-1 ROUGE-2 ROUGE-L METEOR
RSICap Zero-shot $0.03$ $18.31$ $4.33$ $15.21$ $7.35$
GAIA (Synthetic) $1.29$ $24.31$ $5.41$ $18.76$ $15.09$
GIT-10M Zero-shot $0.05$ $17.41$ $3.40$ $13.95$ $6.83$
GAIA (Synthetic) $1.16$ $26.75$ $4.37$ $18.15$ $16.62$
Dataset Mode Text-to-Image Recall (%) Image-to-Text Recall (%)
@5 @10 @20 @50 @100 @5 @10 @20 @50 @100
RSICap Zero-shot 9.79 15.67 23.91 38.99 53.93 10.44 17.02 25.65 39.65 54.16
GAIA (Synthetic) 11.26 17.06 25.80 42.40 58.18 12.83 18.29 26.72 40.68 56.00
GIT-10M Zero-shot 1.83 2.85 4.39 7.66 11.66 2.95 4.54 6.81 11.42 16.60
GAIA (Synthetic) 4.65 6.16 8.38 12.98 17.95 5.12 6.72 9.02 13.53 18.48
TABLE IX: CLIP-ViT-L-14 Classification Performance Comparison between Zero-shot CLIP and RS-domain fine-tuned CLIP models.
Dataset Variant Dataset Size Metrics (%)
Acc1 Acc5 Mean Recall
eurosat CLIP - $62.57$ $95.93$ $63.95$
GAIA (Alt-text) 40K $68.52$ $98.87$ $69.75$
GAIA (Synthetic) 40K $72.46$ $98.91$ $72.68$
SkyCLIP 2.6M $70.54$ $99.41$ $69.88$
RemoteCLIP 165K $56.22$ $93.11$ $55.17$
GeoRSCLIP 5M $74.52$ $99.33$ $75.07$
LAION_RS 726K $71.33$ $99.37$ $72.66$
resisc45 CLIP - $63.35$ $92.40$ $63.82$
GAIA (Alt-text) 40K $65.49$ $91.95$ $66.01$
GAIA (Synthetic) 40K $66.11$ $92.65$ $66.62$
SkyCLIP 2.6M $70.59$ $95.19$ $71.23$
RemoteCLIP 165K $80.57$ $97.87$ $80.82$
GeoRSCLIP 5M $73.59$ $97.89$ $74.19$
LAION_RS 726K $70.11$ $95.18$ $70.63$
imagenet1k CLIP - $75.54$ $94.58$ $75.54$
GAIA (Alt-text) 40K $75.32$ $94.66$ $75.33$
GAIA (Synthetic) 40K $75.49$ $94.57$ $75.48$
SkyCLIP 2.6M $75.56$ $94.68$ $75.55$
RemoteCLIP 165K $64.68$ $88.15$ $64.68$
GeoRSCLIP 5M $66.01$ $90.17$ $66.01$
LAION_RS 726K $75.61$ $94.75$ $75.62$
TABLE X: CLIP-ViT-L-14 Cross-Modal Retrieval Performance Comparison between Zero-shot CLIP and RS-domain fine-tuned CLIP models using GAIA’s thematic tags.
Model Text-to-Image Recall (%) Image-to-Text Recall (%)
@5 @10 @20 @50 @100 @5 @10 @20 @50 @100
CLIP 5.51 9.17 15.09 26.60 38.07 7.41 11.52 17.05 27.55 36.97
GAIA (Alt-text) 5.50 9.67 16.45 29.56 41.78 11.42 17.66 25.52 38.65 49.44
GAIA (Synthetic) 7.76 13.04 20.72 34.28 47.32 12.30 19.07 27.30 40.41 51.79
SkyCLIP 7.67 12.35 18.85 31.26 43.18 6.74 11.06 16.20 25.92 35.80
RemoteCLIP 1.34 2.19 3.30 6.16 9.26 3.84 6.09 9.19 16.76 24.71
GeoRSCLIP 4.36 7.31 10.92 19.30 28.38 5.37 8.82 13.50 22.72 32.13
LAION_RS 7.74 12.99 20.46 33.08 45.03 8.90 13.77 20.07 30.76 41.19